The perimeter of a rectangle equals the sum of the lengths of its four sides if the width of a rectangle is 2 inches and its perimeter is 36 inches, find the length of the rectangle
Answer by lwsshak3(11628)

The perimeter of a rectangle equals the sum of the lengths of its four sides if the width of a rectangle is 2 inches and its perimeter is 36 inches, find the length of the rectangle
**
let x=length of rectangle
let y=width of rectangle
2x+2y=perimeter
2x+2*2=36
2x=36-4=32
x=16
length of rectangle=16